A 1500kg truck slows down from a speed of 25m/s to a complete stop over 3.0s. What is the average force that the brake pads must exert on the truck to complete this motion?
Answer #1

Deceleration =25/3=8.33 m/s²

Mass=1500 kg

Average force by brake pads=8.33*1500=12500 N=12.5 kN